OpenStudy (anonymous):

what is the answer for Q 6q - 1 = -q + 20

OpenStudy (anonymous):

6q + q = 20 + 1 What do you get?

OpenStudy (anonymous):

meh i got 6q + -1 = 20 + -1

OpenStudy (anonymous):

*-1q

OpenStudy (anonymous):

7q = 21 q = 21/7 = ?

OpenStudy (anonymous):

You MUST take all the "q" containing terms to one side, and then do the eqaution.

OpenStudy (anonymous):

okay.. 6q - 1 = -q + 20 6q + -1q = 20 + -1 5q = 19 q = 1.8 ?

OpenStudy (anonymous):

When you take '-q' from the right to the left side, it becomes '+q' or 'q' Numbers change signs when you switch their sides.

OpenStudy (anonymous):

i understand now.. does that also go for the '1' on the left side, or does it inherit the subtraction and bcome negative?

OpenStudy (anonymous):

Yes. ANYTHING, once switches sides, will change its sign.

OpenStudy (anonymous):

so it would be 6q + 1q = 20 + -1 or 6q + 1q = 20 + 1 (: